IMPLEMENTING TWO-COLOR TWO-DIMENSIONAL ELECTRONIC SPECTROSCOPY

\underline{K. L. M. LEWIS}, J. A. MYERS, P. F. TEKAVEC and J. P. OGILVIE, Department of Physics and Biophysics, University of Michigan, Ann Arbor, 48109.

Two-dimensional electronic spectroscopy (2DES) is a powerful tool for revealing electronic coupling and solvation dynamics in condensed phase systems. One of the key problems in implementing 2DES at optical frequencies is maintaining adequate phase stability between the requisite pulse pairs. We present two-color 2DES data obtained with pulse-shaping and diffractive-optics-based approaches and discuss the advantages and disadvantages of the two methods.